OpenStudy (anonymous):

which number is inrational? 81? 169? 156? 144?

OpenStudy (unklerhaukus):

do you mean? √81 √169 √156 √144

OpenStudy (anonymous):

Yes.

OpenStudy (unklerhaukus):

do you know your square numbers? \[1\times1=1^2=1\\2\times2=2^2=4\\3\times3=3^2=9\\4\times4=4^2=16\\5\times5=5^2=25\\\qquad\vdots=\vdots~~=\quad\vdots\]

OpenStudy (anonymous):

Yes Im just making sure. It's a test I really want to pass to bring my grade up.

OpenStudy (unklerhaukus):

the square root of a square number \[\sqrt{n^2}=n\]is than number __ for example \[\sqrt{25}=\sqrt{5^2}=5\]

geerky42 (geerky42):

Yes, so if whether in root is not a square number then it is irrational number.

OpenStudy (anonymous):

Aye. OK. Thank you.

OpenStudy (anonymous):

*Both.

geerky42 (geerky42):

You're welcome.